Shift Summary - OWL Dec 29, 1999

Douglas A. Jensen (djensen@fnal.gov)
Wed, 29 Dec 1999 07:10:03 -0600 (CST)

Shift Summary for OWL shift, 29/Dec/99

On Shift: Bob H, Doug J; Arun for a while.

Hours of beam delivered at all intensities: 7.0/7
Hours of beam delivered at nominal intensity: 6.9
Hours of E799 data (special runs not included): 6.6
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: End/Start run - 20 min, TeV aborts - few spills

========================================================================
Runs ended this shift:
Run 15314 - BEAM_799, 20.7M.
Run 15315 - ongoing - BEAM_799

Tapes brought to FCC this shift: UPA277-9
UPB277-9
UPC277-9
UPS093

UPF068
UPK036
UPM051
========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E799 datataking @ close to 1E13

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15314 00:00 New Shift
45 Mount UPX018
15314 55 Run End "I'm Full!"
Restart SDAQ
15315 1:16 Start Run
1:20 No CsI histograms, follow instructions, get 'em.
2:10 Mount UPE019
3:13 Mount UPH051
Arun leaves, having created more disk space.
5:20 mount split tape UPF069

========================================================================
Follow up Notes on Previous Shift
-SDAQ fresh-started before run 15315

========================================================================
Items to Follow up on Next Shift

-Do a DART fresh start when the run ends.

-keep an eye (ear?) on the fan in rack 161. Replace if down time.

-page Sasha Golossanov if there is a 1+ hour shutdown (not owl shift).

- NOTE 3Y-6 KQ module is not being read out - so monitoring histogram
has a hole in it. This also has an effect on YTF.
This is a known readout problem - Sorry for the late call Jason.....

- Note that the He pressure has not changed much in 3 days ?????????
=========================================================================
While I have your attention: Some historical issues....
When the chilled water system went off on the 25th, the temperature
of the Diffusion pumps went up several degrees. I suspect that the
temperature in the mechanical pumps did the same, though we have no
on-line monitoring of them. So - BEWARE - if the chilled water has
a problem, the Vacuum pumps may be in trouble !

On a slightly older issue: last time the regenerator was moved to the
way out position, the control cable was left connected. I discovered it
because it was trying to operate.. not cool. So I unplugged it.
=========================================================================
On the subject of xerography....
There is the following note posted in the counting room...
"Dear Bob T.
I stopped by KTeV as you requested, but I didn't see any milk and
cookies for me! No new copier for KTeV!
-Kris Kringle "

soooooo another opportunity to solve this long standing problem has
passed us by. Boo Hoo hoo hooo hooo!
========================================================================